40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5267     Entry time: Fri Aug 19 01:46:06 2011
Author: Suresh 
Type: Update 
Category: General 
Subject: IFO alignment 

[Keiko, Jamie, Kiwamu, Anamaria,

We followed the procedure that we laid-out in our elog of yesterday.  We completed the first six steps and we now have the y-arm well aligned to the green beam which passes through the center of of both ETMY and ITMY. 

The IR beam was steered with the PZTs to coincide with the green beam.  The BS was adjusted to see IR beam scatter on a target placed near the center of the ETMX.  And then the AS IR beam was steered to the AS camera by adjusting several components along OM path ( we touched OM1, OM2, OM3, OM4, OM5, OMPO and OM6).  We then looked for IR fringes in the AS port from the Y-arm. But no luck there.  We need to realign the IR beam into the Y-arm cavity axis using the pzts.

We aligned ITMX and PRM to get power recycled Michelson fringes at the AS.

 

ELOG V3.1.3-